OpenStudy (anonymous):

f(x)= ((x-4)(x+5)/((13x+4)(4-x)) What graphical feature occurs at x=4? (select all that apply) hole intercept vertical asymptote

OpenStudy (amistre64):

can you factor out the offending zero?

OpenStudy (anonymous):

what do you mean by "offending zero"?

OpenStudy (anonymous):

ldaniel... 4 is just an intercept when y=0 alongside 5. However, y=0 behaves a like horizontal aymptote which rule the possibility of it being a vertical asymptote out. Hole? yes f(x) is discontinuous there but this is simply because of the horizontal asymptote feature it shows at x=4 when y=0

OpenStudy (anonymous):

so its a hole and a intercept or just a hole?

OpenStudy (anonymous):

its an x-intercept

OpenStudy (amistre64):

its just a hole, you cant intercept something if there is nothing there to intercept

OpenStudy (amistre64):

an offending zero is a value that makes the bottom of a fraction equal zero\[\frac1n\]if n=0, the expression is undefined

OpenStudy (anonymous):

so its just a hole, not a intercept and not a vertical asymptote, right?

OpenStudy (amistre64):

\[f(x)= \frac{(x-4)(x+5)}{(13x+4)(4-x)}\] \[f(x)= \frac{-\cancel{(4-x)}(x+5)}{(13x+4)\cancel{(4-x)}}\] \[f(x)= \frac{-(x+5)}{13x+4}\] we can "remove" the offending zero; and when something is removed, it creates a hole.

OpenStudy (amistre64):

if it cant be removed, it becomes an asymptote

OpenStudy (amistre64):

and, if the value of x doesnt create an offending zero, then its just the value of the function at the stated x ... but in this case, x=4 was offending, and could be removed

OpenStudy (anonymous):

if it could be removed then f(x) is continuous...so where the hole come in?

OpenStudy (amistre64):

the comes in from the fact that the function is undefined at x=4 to begin with; just becasue we can construct an equivalent (but not the same) function has no bearing on the parts that make it up

OpenStudy (amistre64):

we are given f(x) ; we can construct an equivalent g(x) to study the behaviour of f(x) with, but g(x) does not excuse the inherent flaws of f(x)

OpenStudy (amistre64):

think of g(x) as lipstick on a pig :)

OpenStudy (amistre64):

consider it this way; a map gives a representation of a road, but a map is not a road. when a bridge across the road is out, the map still gives us an equivalent view - but no matter what the map looks like, the bridge is still out.